# NAP-Tuning: Neural Augmented Prompt Tuning for Adversarially Robust Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) such as CLIP have demonstrated remarkable capabilities in understanding relationships between visual and textual data through joint embedding spaces. Despite their effectiveness, these models remain vulnerable to adversarial attacks, particularly in the image modality, posing significant security concerns. Building upon our previous work on Adversarial Prompt Tuning (AdvPT), which introduced learnable text prompts to enhance adversarial robustness in VLMs without extensive parameter training, we present a significant extension by introducing the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning). Our key innovations include: (1) extending AdvPT from text-only to multi-modal prompting across both text and visual modalities, (2) expanding from single-layer to multi-layer prompt architectures, and (3) proposing a novel architecture-level redesign through our Neural Augmentor approach, which implements feature purification to directly address the distortions introduced by adversarial attacks in feature space. Our NAP-Tuning approach incorporates token refiners that learn to reconstruct purified features through residual connections, allowing for modality-specific and layer-specific feature correction. Comprehensive experiments demonstrate that NAP-Tuning significantly outperforms existing methods across various datasets and attack types. Notably, our approach shows significant improvements over the strongest baselines under the challenging AutoAttack benchmark, outperforming them by 33.5% on ViT-B16 and 33.0% on ViT-B32 architectures while maintaining competitive clean accuracy. This work establishes the importance of architecture redesign in prompt tuning for adversarial robustness, moving beyond loss-focused approaches to create an adaptive defense mechanism that can identify and rectify adversarial perturbations across embedding spaces.
Index Terms:
Adversarial robustness, Vision-Language models, Prompt tuning, Feature purification, Neural augmentation
@@ I Introduction
While AdvPT demonstrated the potential of prompt tuning for adversarial defense, it exhibited three critical limitations: (1) its restriction to text modality prompts, (2) its reliance on single-layer prompting, and (3) its emphasis on loss function redesign rather than architectural innovations. Following our initial work, several approaches [5, 6] have attempted to address some of these issues, particularly by extending prompting to multiple modalities and layers. However, these incremental improvements fail to address the core architectural challenge of prompt tuning in adversarial settings.
The fundamental limitation across all existing approaches—including both our original AdvPT and subsequent works—is the direct transplantation of prompt tuning techniques from the generalization domain (e.g., CoOp [7] and MaPLe [8]) to adversarial defense without rethinking the underlying architecture. This approach overlooks a critical insight: prompt tuning, originally developed for clean data environments, faces fundamentally different challenges in adversarial scenarios where feature spaces are deliberately corrupted. Adversarial perturbations introduce systematic distortions that cannot be effectively countered through loss modifications alone, regardless of whether the prompting occurs in single or multiple modalities, or at one or multiple layers.
To comprehensively address all three limitations, we propose the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning). Our approach not only extends prompting to multiple modalities and layers but, most critically, fundamentally reconceptualizes prompt architecture for adversarial defense by integrating feature purification mechanisms. The Neural Augmentor framework incorporates token refiners—lightweight neural networks that learn to reconstruct clean feature representations from adversarially perturbed inputs through residual connections. This architecture enables modality-specific and layer-specific feature correction, yielding substantially enhanced adversarial robustness. Fig. LABEL:fig:comparison illustrates the evolution from text-only prompting to our comprehensive NAP-Tuning, demonstrating how our method facilitates clean feature reconstruction from adversarially corrupted inputs across both modalities and multiple network layers.
@@
As an extension of our conference paper [3], this work makes several significant contributions:
  • Multi-modal defense framework: We extend the initial text-only AdvPT to a comprehensive multi-modal framework that incorporates learnable prompts in both text and visual pathways, creating a more effective defense mechanism that addresses vulnerabilities across modalities.
  • Multi-layer prompt architecture: We expand from single-layer prompting to multi-layer prompt architectures, allowing for more nuanced and layer-specific defenses that can address adversarial perturbations manifesting differently at various depths of the network.
  • Neural Augmentor design: Most significantly, we propose a novel architecture-level redesign through our Neural Augmentor approach, which implements feature purification via token refiners. This represents a paradigm shift from previous loss-focused prompt tuning methods to architecture-focused prompt tuning defense, addressing the fundamental limitations of existing approaches.
  • Comprehensive evaluation: We provide extensive experimental validation across multiple datasets and attack types under stricter evaluation settings. Our approach demonstrates significant performance improvements over the previous conference version (AdvPT), achieving average gains of 42.6% under PGD-100 attacks and 47.5% under the more challenging AutoAttack benchmark, establishing new standards for adversarial robustness in VLMs.
@@ II-A Vision-Language Models
VLMs have established themselves as powerful frameworks for bridging visual and textual modalities. These models can be categorized into two principal architectures. The first category comprises generative models built upon Large Language Models (LLMs), such as LLaVA [9] and Qwen-VL [10], which extend text-only LLMs with vision encoders to enable multimodal understanding and generation.
The second category—which is the focus of adversarial prompt tuning techniques—utilizes contrastive learning to establish joint embeddings of images and text. Models such as CLIP [1] and ALIGN [2] project both modalities into a unified semantic space where related content exhibits higher similarity. These models demonstrate exceptional capabilities across diverse tasks through their dual-encoder architecture. However, their susceptibility to adversarial attacks, particularly in the visual domain, poses significant security concerns as their deployment expands into critical applications.
@@ II-B Prompt Learning
Prompt learning emerged in natural language processing as an efficient parameter-efficient adaptation technique for pre-trained language models [11, 12]. This paradigm focuses on optimizing input transformations rather than modifying model parameters.
For VLMs, text-based prompt learning methods such as CoOp [7] and CoCoOp [13] demonstrated that replacing hand-crafted templates with learnable context vectors could significantly improve model performance. Our previous work, AdvPT [3], applied this concept to adversarial defense, showing that learnable text prompts could enhance robustness against adversarial attacks without modifying the underlying model. Multi-modal prompt learning approaches including MaPLe [8] and PromptKD [14] subsequently extended this paradigm by incorporating prompts in both textual and visual pathways. These architectural innovations have provided foundational techniques for multi-modal adversarial prompt tuning, though their primary focus on generalization rather than robustness presents limitations when applied to adversarial defense.
@@ II-C Adversarial Prompt Tuning
Conventional adversarial defense methods such as adversarial training [4], TRADES [15], and MART [16] have proven effective but remain computationally prohibitive for large-scale VLMs. These approaches typically require training the whole model parameters from scratch, making them impractical for deployment in VLMs. An innovative alternative approach was introduced by Mao et al. [17] through their TeCoA framework, which intelligently leverages pre-trained VLM weights as initialization and selectively finetunes CLIP’s vision encoder.
Our previous work, AdvPT [3], introduced a more efficient and more effective paradigm by applying prompt learning to adversarial defense in VLMs. By aligning text embeddings with adversarial image embeddings through learnable prompts, AdvPT demonstrated that robust representations could be achieved with minimal computational overhead. Recent extensions to this approach include FAP [5], which leverages limited data to learn adversarially correlated text supervision while enhancing multi-modal feature consistency and uni-modal feature differentiation between natural and adversarial examples.
Recent advances in test-time optimization have yielded particularly impressive results in defending VLMs against adversarial attacks. Sheng et al. [18] developed R-TPT, an elegant test-time prompt tuning method that enhances the robustness of vision-language models by minimizing pointwise entropy and introducing a reliability-based weighted ensembling strategy. Their approach demonstrates the power of adaptive defenses that operate at inference time. TAPT [19] introduces an unsupervised approach that optimizes bimodal prompts at test time to enhance CLIP’s robustness while maintaining clean performance. In parallel, Xing et al. [20] introduced TTC, an innovative training-free defense mechanism that ingeniously leverages CLIP’s pre-trained vision encoder to counteract adversarial images during inference. Their approach offers complementary protection that can be effectively combined with existing robust models, demonstrating the value of ensemble strategies in comprehensive defense frameworks.
While these methods have achieved incremental improvements through loss function innovations, they overlook a critical consideration: the fundamental mismatch between prompt architectures designed for generalization and the unique requirements of adversarial robustness. Addressing adversarial vulnerabilities requires mechanisms specifically engineered to counteract feature-level distortions—a challenge that far exceeds general-purpose representation enhancement. This insight motivates our current work, which demonstrates that architectural innovations in prompt design can substantially improve adversarial robustness beyond what is possible through loss function modifications alone. By reimagining prompt architecture specifically for adversarial defense, we address the feature distortion mechanisms that underlie successful attacks on VLMs.
@@ III Preliminary: Adversarial Prompt Tuning
To provide context for our extended approach, we first revisit the key concepts of Adversarial Prompt Tuning (AdvPT) as introduced in our previous work [3].
@@ III-B Adversarial Attacks on VLMs
$x_{adv}$ $=\arg\max_{x^{\prime}}\mathcal{L}_{CE}(f_{I}(x^{\prime}),f_{T}(t_{y}))$
@@
where $t_{y}$ is the text template for class $y$, and $\mathcal{L}_{CE}$ is the cross-entropy loss measuring image-text similarity.
@@ IV Neural Augmentor for Multi-Modal Adversarial Prompt Tuning
Adversarial attacks fundamentally operate by introducing carefully crafted perturbations to input data that cause neural networks to make incorrect predictions. In the context of VLMs, these perturbations manipulate feature representations in ways that disrupt the crucial alignment between visual and textual information. Building upon our preliminary work on AdvPT, we introduce a comprehensive redesign of prompt-based adversarial defenses through our NAP-Tuning framework. This section details our approach, its architectural innovations, and theoretical foundations.
@@ IV-A Framework Overview
Fig. LABEL:fig:comparison (c) introduces the overall framework, while Fig. 3 illustrates the architecture of our Neural Augmentor module. At its core, our approach aims to defend VLMs against adversarial attacks by addressing their fundamental vulnerability: the distortion of feature representations. Unlike previous prompt defenses that focus on loss function modifications or robust classification boundaries, our method directly targets the feature corruption mechanism through specialized architectural components.
The framework consists of three primary components that work in concert. First, a multi-modal prompting system introduces learnable prompt vectors that operate simultaneously in both textual and visual pathways of the model. Second, a multi-layer prompt architecture implements hierarchical prompting structures at different depths within the transformer architecture to address adversarial effects at various levels of feature abstraction. Third, Neural Augmentor modules serve as specialized neural networks that perform feature purification through token refiners.
@@ IV-C1 Feature Distortion Theory
where $\Phi$ is a learned correction function that approximates $-\Delta f$, effectively “purifying” the adversarial features (including both intermediate and final feature representations) back toward their clean counterparts.
@@ IV-C2 TokenRefiner Architecture
The core component of the Neural Augmentor is the TokenRefiner $R$—a lightweight neural network that processes individual token representations to identify and correct adversarial distortions. For each token representation $z\in\mathbb{R}^{d}$, the TokenRefiner computes a corrective term $\tilde{z}=R(z)$. The TokenRefiner function is implemented as a two-layer network with a residual connection. It ensures stable gradient flow during training, facilitating effective optimization. Moreover, when no correction is needed (e.g., for clean inputs), the network can learn to output values near zero, effectively preserving the original features through an identity fallback mechanism. The $R$ operates on both text and visual tokens, with modality-specific parameters that allow it to learn distinct correction patterns for each pathway.
@@
The Neural Augmentor implements a feature purification mechanism that operates across three primary dimensions. At the token level, each representation is individually refined, allowing for localized corrections that address token-specific perturbations. At the layer level, TokenRefiner deployed at different network depths learn to address distinct types of perturbations, from low-level feature distortions to high-level semantic shifts. At the modality level, separate TokenRefiner for text and visual pathways allow for specialized correction patterns that address the unique vulnerabilities of each modality.
This multi-dimensional purification approach enables a comprehensive defense against diverse adversarial attacks, addressing perturbations at their source—the feature representations themselves. Rather than treating adversarial robustness as a classification boundary problem, our approach directly targets the mechanism by which adversarial examples cause misclassifications: the distortion of feature representations. By restoring these representations to their clean counterparts, we enable the model to maintain its performance even in the presence of adversarial perturbations.
@@ IV-D2 Principled Loss Formulation
TABLE I: Evaluation results on ViT-B16 under clean and adversarial settings. Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (CWA and AGS), where $\dagger$ presents our method. Best results are highlighted in bold.
@@
FAP $59.7$ $92.2$ $60.8$ $67.8$ $25.4$ $71.3$ $84.5$ $87.1$ $61.1$ $64.0$ $75.3$ $70.6$
AdvPT $67.5$ $94.1$ $69.6$ $69.0$ $28.7$ $85.0$ $87.9$ $91.7$ $70.8$ $70.7$ $77.4$ $73.8$
AdvPT-V $60.7$ $89.8$ $36.9$ $28.1$ $16.1$ $63.8$ $54.4$ $84.7$ $51.7$ $58.4$ $54.1$ $57.0$
AdvPT-VLI $63.8$ $85.7$ $37.2$ $18.2$ $12.0$ $62.3$ $53.6$ $78.7$ $46.9$ $52.8$ $50.5$ $55.4$
AdvPT-VLJ $65.9$ $88.0$ $37.8$ $20.0$ $10.3$ $69.6$ $59.7$ $82.1$ $53.5$ $56.6$ $55.8$ $60.0$
@@
CWA Vanilla $50.1$ $91.5$ $40.5$ $18.3$ $21.9$ $76.1$ $66.7$ $85.9$ $60.9$ $59.8$ $61.8$ $61.1$
FAP $49.6$ $91.8$ $56.8$ $23.2$ $24.3$ $66.6$ $82.2$ $85.4$ $58.6$ $61.2$ $70.6$ $65.6$
AdvPT $52.1$ $93.1$ $61.8$ $46.9$ $30.8$ $74.9$ $88.2$ $89.8$ $67.8$ $70.5$ $74.3$ $73.0$
AdvPT-V $51.2$ $88.4$ $36.0$ $23.5$ $15.4$ $56.5$ $53.3$ $81.9$ $48.3$ $55.8$ $51.6$ $51.3$
AdvPT-VLI $54.4$ $84.7$ $35.4$ $17.7$ $11.4$ $55.5$ $52.5$ $76.9$ $45.2$ $50.1$ $48.5$ $50.8$
AdvPT-VLJ $53.0$ $86.5$ $35.2$ $16.4$ $8.5$ $52.6$ $50.7$ $77.7$ $42.5$ $51.0$ $48.6$ $49.1$
@@
AGS Vanilla $53.5$ $86.8$ $30.9$ $11.0$ $18.1$ $56.9$ $56.9$ $78.0$ $54.1$ $53.7$ $49.7$ $54.3$
FAP $51.1$ $88.2$ $50.2$ $12.3$ $22.2$ $58.9$ $77.4$ $80.1$ $53.4$ $55.0$ $61.3$ $58.5$
AdvPT $54.6$ $88.5$ $51.5$ $26.1$ $26.1$ $55.9$ $76.0$ $80.4$ $62.6$ $62.8$ $64.1$ $61.5$
AdvPT-V $53.0$ $85.8$ $33.2$ $15.6$ $13.6$ $47.6$ $47.7$ $78.2$ $44.3$ $51.3$ $44.3$ $50.8$
AdvPT-VLI $56.3$ $81.1$ $33.6$ $14.7$ $10.6$ $46.2$ $48.5$ $74.3$ $41.5$ $46.2$ $41.7$ $47.5$
AdvPT-VLJ $54.9$ $83.0$ $31.6$ $12.1$ $7.5$ $44.0$ $45.3$ $73.9$ $38.6$ $46.2$ $41.4$ $47.3$
[…]
PGD Vanilla $2.7$ $28.3$ $2.0$ $0.1$ $0.0$ $14.4$ $3.2$ $8.8$ $1.4$ $1.4$ $4.0$ $6.7$
FAP $24.0$ $62.4$ $26.7$ $0.0$ $5.2$ $21.8$ $51.7$ $33.7$ $19.5$ $25.4$ $10.9$ $23.3$
AdvPT $1.5$ $27.0$ $6.6$ $0.2$ $0.6$ $0.9$ $4.3$ $2.8$ $0.7$ $1.6$ $2.1$ $5.2$
AdvPT-V $19.6$ $61.5$ $18.6$ $8.1$ $3.8$ $12.4$ $25.0$ $38.0$ $9.8$ $17.3$ $17.0$ $20.4$
AdvPT-VLI $21.8$ $59.1$ $18.6$ $10.0$ $2.9$ $12.9$ $21.7$ $38.2$ $10.0$ $15.7$ $14.2$ $20.2$
AdvPT-VLJ $20.0$ $58.7$ $16.4$ $10.2$ $2.3$ $11.4$ $20.5$ $32.8$ $8.6$ $14.9$ $12.7$ $18.9$
@@
AdvPT-V $14.9$ $55.3$ $15.0$ $2.2$ $2.2$ $8.6$ $18.8$ $32.1$ $5.9$ $12.5$ $13.0$ $17.9$
AdvPT-VLI $17.0$ $54.3$ $14.3$ $9.2$ $1.6$ $9.1$ $16.2$ $33.6$ $4.8$ $11.4$ $11.3$ $17.0$
AdvPT-VLJ $10.3$ $43.9$ $10.3$ $3.7$ $1.2$ $4.4$ $9.7$ $19.8$ $1.9$ $6.7$ $6.0$ $11.1$
@@
TABLE II: Evaluation results on ViT-B32 under clean and adversarial settings. Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (CWA and AGS), where $\dagger$ presents our method. Best results are highlighted in bold.

### v2

@@ Abstract
Vision-Language Models (VLMs) such as CLIP have demonstrated remarkable capabilities in understanding relationships between visual and textual data through joint embedding spaces. Despite their effectiveness, these models remain vulnerable to adversarial attacks, particularly in the image modality, posing significant security concerns. Building upon our previous work on Adversarial Prompt Tuning (AdvPT), which introduced learnable text prompts to enhance adversarial robustness in VLMs without extensive parameter training, we present a significant extension by introducing the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning). As a significant extension, NAP-Tuning first establishes a comprehensive multi-modal (text and visual) and multi-layer prompting framework. The core of this framework is a targeted structural augmentation for feature-level purification, implemented through our Neural Augmentor approach. This framework implements feature purification by incorporating TokenRefiners—lightweight neural modules that learn to reconstruct purified features via residual connections—to directly address distortions in the feature space. This structural intervention is what enables the multi-modal and multi-layer system to effectively perform modality-specific and layer-specific feature rectification. Comprehensive experiments demonstrate that NAP-Tuning significantly outperforms existing methods across various datasets and attack types. Notably, our approach shows significant improvements over the strongest baselines under the challenging AutoAttack benchmark, outperforming them by 32.3% on ViT-B16 and 31.3% on ViT-B32 architectures while maintaining competitive clean accuracy. This work highlights the efficacy of internal feature-level intervention in prompt tuning for adversarial robustness, moving beyond input-side alignment approaches to create an adaptive defense mechanism that can identify and rectify adversarial perturbations across embedding spaces.
@@ I Introduction
While AdvPT demonstrated the potential of prompt tuning for adversarial defense, it exhibited three critical limitations: (1) its restriction to text modality prompts, (2) its reliance on single-layer prompting, and (3) its emphasis on loss function redesign rather than architectural innovations. Following our initial work, several approaches [41, 15] have attempted to address some of these issues, particularly by extending prompting to multiple modalities and layers. However, these works, while incrementally beneficial, share the same fundamental architectural limitations as AdvPT and thus fail to address the core challenge of robust feature representation.
The fundamental limitation across all existing prompt-based approaches is the direct transplantation of prompt tuning techniques from the generalization domain (e.g., CoOp [40] and MaPLe [8]) to adversarial defense without rethinking the underlying architecture. This overlooks a critical insight: adversarial robustness is a fundamentally harder task than standard generalization, as it requires richer model capacity [25, 36]. This requirement for increased capacity creates a fundamental trade-off during VLM adaptation. While freezing the backbone preserves pre-trained knowledge, the limited structural capacity of standard prompt tuning, which typically optimizes a small set of input-side vectors, often fails to rectify corrupted features effectively. Conversely, high-capacity full fine-tuning risks catastrophic forgetting of pre-trained representations [28]. We aim to bridge this gap by incorporating a modular structural augmentation that provides targeted capacity for feature purification while maintaining the backbone’s integrity. A systematic analysis and positioning of these adaptation paradigms are further detailed in Section II-D.
To address these limitations, we propose the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning). Our approach introduces a targeted structural augmentation designed to navigate the trade-off between model capacity and knowledge preservation. We keep the backbone frozen to avoid catastrophic forgetting, but introduce a modular neural component: the Neural Augmentor. Unlike standard prompt vectors, the Augmentor is a set of lightweight neural networks (TokenRefiners) that operate directly on the feature representations. This modular design provides the functionality required to model and rectify feature distortions at the token level. Fig. 3(a) illustrates the evolution from text-only prompting to our comprehensive NAP-Tuning.
(a) Comparison of adversarial prompt tuning approaches: (a) The original AdvPT that uses only text prompts, (b) Existing multimodal approaches that incorporate prompts in both visual and text pathways, and (c) Our proposed NAP-Tuning framework that extends multimodal prompting with feature purification via token refiners, enabling the reconstruction of clean feature representations from adversarially perturbed inputs.
@@
As an extension of our conference paper [37], which introduced the concept of Adversarial Prompt Tuning, this work makes several significant contributions:
  • Systematic Extension: Building upon our original text-only AdvPT [37], we implement a coordinated framework across both visual and textual pathways. This systemic extension addresses adversarial vulnerabilities that manifest across modalities, providing a unified multi-layer prompting architecture for robust vision-language adaptation.
  • Modular Structural Augmentation: We introduce the Neural Augmentor, a modular structural augmentation utilizing lightweight TokenRefiners for explicit feature-level purification. This design represents a functional shift from loss-alignment methods to direct feature rectification, providing a plug-and-play mechanism that is easily compatible with various frozen VLM backbones.
  • Comprehensive Validation: We provide extensive experimental validation for our method across 11 datasets and under stricter evaluation settings. Our analysis confirms the effectiveness of the feature purification mechanism, demonstrating state-of-the-art robustness gains while preserving generalization.
TABLE I: Comparative analysis of Adversarial Adaptation Paradigms for VLMs. We analyze different strategies based on their architectural approach to the robustness-generalization trade-off.
Adaptation Paradigm (Examples) Preserves Prior Increases Model Architectural Trainable Resulting Robustness
Knowledge? Capacity? Redesign? Parameters Potential
(No Adaptation)
Full Fine-Tuning (TeCoA [18]) N/A a High Low
Prompt Tuning (FAP [41], APD [15]) Low Low Medium
Neural Augmentation (Ours) High Low-Medium High
  • a
    This paradigm does not add new capacity, but unlocks all existing (Very High) parameters for training.
@@ II-A Vision-Language Models
VLMs establish a semantic bridge between visual and textual modalities, typically following two architectural paradigms: generative models based on Large Language Models (e.g., LLaVA [13], Qwen-VL [35]) and contrastive models (e.g., CLIP [23], ALIGN [7]) that learn joint embedding spaces. While the latter demonstrates exceptional zero-shot transfer capabilities through unified multimodal projections, their inherent susceptibility to adversarial perturbations—particularly in the visual domain—poses significant security risks. Ensuring the robustness of these contrastive embeddings is critical as their deployment expands into safety-critical applications.
@@ II-B Prompt Learning
Originating as a parameter-efficient alternative to full fine-tuning in NLP [14, 10], prompt learning adapts pre-trained models by optimizing input-side transformations. In the VLM context, methods such as CoOp [40] and CoCoOp [39] replaced manual templates with learnable context vectors. Subsequent innovations, including MaPLe [8] and PromptKD [12], extended this paradigm to multi-modal pathways. However, as these techniques were primarily engineered for task generalization, their fixed-architecture design presents a fundamental bottleneck when attempting to counteract the structural feature distortions introduced by adversarial attacks.
@@ II-C Adversarial Prompt Tuning
To circumvent the prohibitive computational costs of conventional adversarial training [16, 36, 31], AdvPT [37] introduced an efficient defense paradigm by aligning text prompts with adversarial image embeddings. This prompted several extensions, such as FAP [41], which focuses on few-shot adversarial consistency, and APD [15], which leverages bimodal distillation for robust adaptation. In parallel, test-time optimization strategies (e.g., R-TPT [26], TAPT [30], TTC [34]) have enhanced VLM security during inference. Despite these advancements, existing methods predominantly focus on loss-level alignment of input-side prompt vectors. Our work diverges from these trends by introducing a modular structural augmentation designed specifically for internal feature-level rectification, a mechanism that targets the root cause of feature distortion beyond simple distribution alignment.
II-D Architectural Trade-offs in Robust VLM Adaptation
The theoretical foundation for adapting VLMs for robustness involves navigating the fundamental trade-off between standard accuracy and adversarial robustness [36]. Robust learning theory establishes that achieving adversarial generalization is a more difficult task, requiring significantly higher model capacity than standard generalization [25]. In the context of large pre-trained models, this requirement creates a specific dilemma: while full fine-tuning provides the necessary capacity to learn robust features, it risks the degradation of pre-trained knowledge [28, 38]. Conversely, fixed-architecture prompt tuning preserves generalization but lacks the structural complexity needed to counteract feature-level distortions.
We clarify this landscape in Table I, positioning our work within three dominant adaptation paradigms. Unlike conventional prompt tuning that relies on loss-alignment of input-side vectors, our Neural Augmentation paradigm introduces a modular structural intervention. By incorporating internal TokenRefiners, we provide the targeted capacity required for feature purification without compromising the backbone’s integrity. This design is fundamentally modular, allowing the rectification mechanism to be integrated into various frozen transformer backbones as a generic plug-in.
@@ III Preliminary: Adversarial Prompt Tuning
To provide context for our extended approach, we first revisit the key concepts of Adversarial Prompt Tuning as introduced in our previous work [37].
@@ III-B Adversarial Attacks on VLMs
$x_{adv}$ $=\arg\max_{x^{\prime}}\mathcal{L}_{CE}(cos(f_{I}(x^{\prime}),f_{T}(\mathbf{T})),y)$
@@
where $\mathbf{T}=\{t_{1},...,t_{k}\}$ is the set of all $k$ class text templates, $y$ is the ground-truth class label, $cos(\cdot,\cdot)$ is the cosine similarity function which produces the logits, and $\mathcal{L}_{CE}$ is the standard cross-entropy loss.
@@ IV Neural Augmentor for Multi-Modal Adversarial Prompt Tuning
Building upon our preliminary work on AdvPT, we introduce the NAP-Tuning framework as a structural augmentation for prompt-based adversarial defenses. This section details our approach, its structural augmentations, and theoretical foundations.
@@ IV-A Framework Overview
Fig. 3(a) (c) and Fig. 3 illustrate the NAP-Tuning framework and its core Neural Augmentor module. Conceptually, our approach represents a strategic evolution of the efficiency-oriented paradigm established by AdvPT [37]. While the original AdvPT focused on minimizing computational overhead by avoiding backbone backpropagation, NAP-Tuning identifies a superior trade-off: by incorporating lightweight structural augmentations for internal feature purification, we achieve a massive leap in robustness while maintaining an efficiency profile that remains orders of magnitude superior to traditional full fine-tuning. This conceptual shift—moving from purely input-side alignment to internal feature rectification—is implemented through three coordinated components: (1) a multi-modal prompting system for visual-textual coordination, (2) a multi-layer prompt architecture for hierarchical intervention, and (3) Neural Augmentor modules that perform explicit token-level purification via residual refiners.
@@ IV-C1 Feature Distortion Theory
where $\Phi$ is a learned correction function that approximates $\Delta f$, effectively “purifying” the adversarial features (including both intermediate and final feature representations) back toward their clean counterparts.
@@ IV-C2 TokenRefiner Architecture
The core component of the Neural Augmentor is the TokenRefiner RR—a lightweight neural network that processes individual token representations to identify and correct adversarial distortions. For each token representation $z\in\mathbb{R}^{d}$, the TokenRefiner computes a corrective term $\tilde{z}=R(z)$. The TokenRefiner function is implemented as a two-layer network with a residual connection. It ensures stable gradient flow during training, facilitating effective optimization. Moreover, when no correction is needed (e.g., for clean inputs), the network can learn to output values near zero, effectively preserving the original features through an identity fallback mechanism. The $R$ operates on both text and visual tokens, with modality-specific parameters that allow it to learn distinct correction patterns for each pathway.
@@
The Neural Augmentor implements a hierarchical purification mechanism across token, layer, and modality dimensions, enabling localized and structural rectification of adversarial artifacts. This approach represents a conceptual shift from treating robustness as a classification boundary problem (typical of input-side alignment methods) to directly targeting the root cause: the distortion of internal feature representations. By explicitly intercepting and restoring perturbed tokens within the backbone’s feature flow, our design maintains the model’s representational integrity and cross-modal alignment even under sophisticated adversarial threats.
@@ IV-D2 Principled Loss Formulation
TABLE II: Evaluation results on ViT-B16 under clean and adversarial settings. Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (M-Attack), where $\dagger$ presents our method. Best results are highlighted in bold.
@@
FAP $60.1$ $92.8$ $60.7$ $68.3$ $26.0$ $72.5$ $84.6$ $87.7$ $60.9$ $65.2$ $75.1$ $68.5$
APD $57.7$ $90.2$ $57.3$ $65.7$ $28.9$ $78.2$ $85.7$ $84.7$ $58.8$ $59.6$ $75.8$ $67.5$
AdvPT 67.5 94.1 69.6 $69.0$ $28.7$ $85.0$ $87.9$ 91.7 $70.8$ 70.7 77.4 73.9
AdvPT-V $60.7$ $89.8$ $36.9$ $28.1$ $16.1$ $63.8$ $54.4$ $84.7$ $51.7$ $58.4$ $54.1$ $54.4$
AdvPT-VLI $63.8$ $85.7$ $37.2$ $18.2$ $12.0$ $62.3$ $53.6$ $78.7$ $46.9$ $52.8$ $50.5$ $51.1$
AdvPT-VLJ $65.9$ $88.0$ $37.8$ $20.0$ $10.3$ $69.6$ $59.7$ $82.1$ $53.5$ $56.6$ $55.8$ $54.5$
@@
M-Attack Vanilla $21.3$ $37.3$ $14.6$ $16.8$ $13.9$ $35.2$ $21.5$ $38.6$ $20.1$ $33.0$ $26.9$ $25.4$
FAP $43.6$ $70.5$ $30.1$ $11.6$ $24.0$ $48.2$ $42.5$ $55.7$ $42.6$ $45.5$ $36.8$ $41.0$
APD $41.9$ $71.6$ $25.4$ $15.7$ $19.0$ $42.3$ $36.4$ $57.0$ $41.4$ $40.3$ $33.3$ $38.6$
AdvPT $34.4$ $43.6$ $12.8$ $8.8$ $9.6$ $38.5$ $29.0$ $47.5$ $32.6$ $39.0$ $45.8$ $31.1$
AdvPT-V $45.3$ $67.3$ $25.7$ $13.6$ $17.8$ $46.3$ $39.5$ $56.4$ $44.4$ $39.7$ $37.4$ $39.4$
AdvPT-VLI $47.9$ $66.4$ $31.5$ $14.8$ $16.1$ $46.7$ $40.6$ $53.2$ $49.7$ $48.0$ $45.7$ $41.9$
AdvPT-VLJ $40.6$ $71.7$ $24.8$ $18.9$ $19.5$ $45.4$ $40.2$ $51.9$ $48.5$ $50.2$ $54.7$ $42.4$
@@
PGD Vanilla $2.7$ $28.3$ $2.0$ $0.1$ $0.0$ $14.4$ $3.2$ $8.8$ $1.4$ $1.4$ $4.0$ $6.0$
FAP $22.3$ $67.2$ $25.5$ $30.5$ $22.9$ $20.7$ $75.3$ $46.6$ $33.0$ $31.9$ $41.1$ $37.8$
APD $23.6$ $71.4$ $31.5$ $31.0$ $23.9$ $25.2$ $74.3$ $44.9$ $48.7$ $33.3$ $40.0$ $40.7$
AdvPT $1.5$ $27.0$ $6.6$ $0.2$ $0.6$ $0.9$ $4.3$ $2.8$ $0.7$ $1.6$ $2.1$ $4.4$
AdvPT-V $19.6$ $61.5$ $18.6$ $8.1$ $3.8$ $12.4$ $25.0$ $38.0$ $9.8$ $17.3$ $17.0$ $21.0$
AdvPT-VLI $21.8$ $59.1$ $18.6$ $10.0$ $2.9$ $12.9$ $21.7$ $38.2$ $10.0$ $15.7$ $14.2$ $20.5$
AdvPT-VLJ $20.0$ $58.7$ $16.4$ $10.2$ $2.3$ $11.4$ $20.5$ $32.8$ $8.6$ $14.9$ $12.7$ $19.0$
@@
AdvPT-V $14.9$ $55.3$ $15.0$ $2.2$ $2.2$ $8.6$ $18.8$ $32.1$ $5.9$ $12.5$ $13.0$ $16.4$
AdvPT-VLI $17.0$ $54.3$ $14.3$ $9.2$ $1.6$ $9.1$ $16.2$ $33.6$ $4.8$ $11.4$ $11.3$ $16.6$
AdvPT-VLJ $10.3$ $43.9$ $10.3$ $3.7$ $1.2$ $4.4$ $9.7$ $19.8$ $1.9$ $6.7$ $6.0$ $10.7$
@@
TABLE III: Evaluation results on ViT-B32 under clean and adversarial settings. Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (M-Attack), where $\dagger$ presents our method. Best results are highlighted in bold.